OpenStudy (anonymous):

solve the equation 1/2x + 3/4 = 1/4 for x

OpenStudy (texaschic101):

there is 2 ways to do this...one with fractions and one without fractions with : 1/2x + 3/4 = 1/4 --- subtract 3/4 from both sides 1/2x = 1/4 - 3/4 1/2x = - 1/2 -- divide both sides by 1/2 x = (-1/2) / (1/2) x = -1/2 * 2 x = -1 without : 1/2x + 3/4 = 1/4 --- multiply everything by common denominator of 4 2x + 3 = 1 --- subtract 3 from both sides 2x = -2 -- divide by 2 x = -1 either way you should arrive at the same answer....you just need to decide on which way is easier for you to do

OpenStudy (anonymous):

Okay thank you!

OpenStudy (texaschic101):

sure thing :)